# Anticipatory Music Transformer

### v1

@@ 7 Ethical Considerations for Generative Models of Music
We find it likely that generative models of music will become widely deployed and used within the next several years (if not sooner). We are particularly concerned about the economic implications of this technology: how will these models affect labor markets for creative work (Section 7.1)? We are uncertain about the legal status of these models: how will music created using models be treated by intellectual property and copyright law (Section 7.2)? We are also curious how the advent of these models will change our perceptions and experience of art and creativity. Will widespread deployment of models trained on Western music reinforce existing Western hegemony in music culture (Section 7.3)? Could the use of these models lead to creative stagnation, or loss of interest in music as an art form (Section 7.4)? Analogous questions and concerns have been raised in the broader context of research on generative models, but the focus of this discussion is often grounded in generation of language or images. These discussions are worth revisiting in the specific context of music generation.
7.1 The Creative Economy and Labor Markets
@@
Music composition has economic value. What happens to the music economy—and to the people who rely on it for their livelihoods—if its creative product can be automatically generated by a machine? Economic research suggests that new technologies exert two countervailing effects on the labor market: a displacement effect whereby automation reduces demand for labor, and a productivity effect whereby greater productivity increases demand for labor (Acemoglu and Restrepo, 2018). It is unclear to us what the net effect of these two forces will be on the size of the labor market for music composition. If latent demand for new music is high, then decreasing the cost of music production could lead to more employment opportunities for composers. That said, this reconfiguration of the music economy might be disruptive, and success in the new music economy could require a new skillset that is difficult (or unappealing) for current workers to acquire. We suspect that generative music technology will affect the current music economy, and may displace some workers.
@@ 7.2 Copyright
Regarding the outputs of a generative music model: generative models have been shown to plagiarize training data in both the language (Carlini et al., 2022) and vision (Somepalli et al., 2022) domains. Based on this evidence, we are certain that generative models of music are similarly capable of plagiarism. Technical mitigation strategies to prevent this behavior based on differential privacy or near access-freeness are nascent (Vyas et al., 2023), as are socio-technical strategies, such as providing copyright holders an opt-out mechanism to exclude their content from training datasets (Henderson et al., 2023). The models trained in this paper provide no technical guarantees against plagiarism. Therefore, any outputs of this model should be presumed to risk infringing on copyrighted material.
@@
7.3 Western Bias
While the methods developed in this paper are broadly applicable to modeling diverse musical traditions, the MIDI data format and the Lakh MIDI dataset used to train the anticipatory music transformer both impose a strong bias towards the Western music tradition. In Section 2 we adopted a musical vocabulary that describes musical pitch according to the 12-tone chromatic scale, which we inherit from the MIDI format of our training data. Models trained using this vocabulary are incapable of expressing music outside the 12-tone scale. This excludes, for example, the Koron and Sori quarter tones of Persian Dastgah music (Abdoli, 2011), the 53 Koma of Turkish makam (Benetos and Holzapfel, 2013), and the continuous-pitch gamaka appearing in Carnatic music (Viraraghavan et al., 2018). For further discussion of the consequences of modeling using a 12-tone discrete scale, see Lenchitz (2021).
@@ 7.4 Models and Art
7.5 Releasing the Anticipatory Music Transformer
@@
Acknowledgements
This work was done at the Center for Research on Foundation Models (CRFM). We thank the CRFM and the Stanford Institute for Human-Centered Artificial Intelligence (HAI) for supporting this work. Toyota Research Institute (TRI) and Schmidt Futures provided funds to support this work. Most of the experiments discussed in this paper were conducted on Cloud TPU VMs, provided by Google’s TPU Research Cloud (TRC).
Appendix A Licensing
@@
Appendix B Encoding Details for Anticipatory Infilling Models.
The tokenization of training examples (Section 3.3) for anticipatory infilling models using an arrival-time encoding of events (Definition 2.2) is described by Definition B.1. As discussed in Section 3.5, we double the base vocabulary to distinguish between anticipated events and non-anticipated events. We also include a sequence separator token and global control codes $\mathbf{z}\in\{\texttt{AR},\texttt{AAR}\}$ (Section 3.3) as well as the REST token (Section 3.2).
Definition B.1.
@@
Example B.2.
See Table 4 for a structured description of this sequence.
The tokenization of training sequences for autoregressive models using an interarrival-time encoding (Defintion 2.3) is described by Definition B.3. We double the base vocabulary of notes and instruments to distinguish between onsets and offsets. We truncate interarrival times longer than 10 seconds (to 10 seconds).
Definition B.3.
[…]
If $\mathbf{x}^{\prime}_{i}$ is an offset,
[…]
And regardless,
[…]
Example B.4.
@@ 00 $10,048$ $11,060$ 00s $480$ms C4 piano
Appendix C A Prior over Music Infilling Controls
@@
Augmenting by a factor of $30$ results in an arrival-time encoded Lakh MIDI train set of approximately $52$B tokens. For training examples of length $1024$ and batch size $512$, this results in approximately one epoch per $100,000$ optimization steps. For models trained with larger step counts, we take multiple passes over the augmented dataset. Given the relatively small size of the Lakh MIDI dataset, anticipation may be of interest as a regularization technique (see Appendix J for some evidence of possible regularizing effects attributable to anticipation). In this case, it could be fruitful to (i) increase the augmentation factor further during training, and (ii) place a smaller weight on training without anticipation, to minimize the amount of duplication of the original (unaugmented) dataset within the augmented dataset.
Appendix D Details of the Lakh MIDI Dataset
@@
Figure 3: The distribution of instantaneous tokens/second calculated for the arrival-time tokenized Lakh MIDI validation split. Mean instantaneous tokens/second for the Lakh MIDI dataset is $68$ with a standard deviation of $51$ events/second.
Appendix E Measuring Cross-Entropy in Bits per Second
Because the arrival-time tokenization (Definition 2.2) and the interarrival-time tokenization (Defintion 2.3) describe nearly equivalent information, the log-loss of models trained using either encoding can be meaningfully compared via a unit conversion. To make comparison agnostic to encoding, we report losses in units of bits per second: this is the total log-loss of the test set, divided by the number of seconds of music in the test set. Concretely, given a per-token loss $L$ reported in nats per token, conversions for interrarival time and arrival time encodings of the test set defined in Section D are:
@@ Appendix E Measuring Cross-Entropy in Bits per Second
Truncations during tokenization (see Section 2.1) result in slightly different information content between an arrival-time tokenized sequence and the corresponding interarrival-time tokenized sequence. In particular, $0.05$% of note durations in the Lakh MIDI dataset exceed $10$ seconds, and are truncated (to $10$ seconds) under arrival-time tokenization. Likewise, a total of $3947$ interarrival times ($0.00$%) in the Lakh MIDI dataset exceed $10$ seconds, and are truncated (to $10$ seconds) under interarrival-time tokenization. Finally, $0.09$% of training examples under arrival-time tokenization exceed $100$ seconds in length, and are discarded from the arrival-time tokenized dataset. These instances are sufficiently rare that they exert a negligible effect on the information content (and therefore on bits per second comparisons) between the arrival-time and interarrival-time tokenized sequences.
Appendix F Hyperparameters and Resources
@@
 Optimizer AdamW [Loshchilov and Hutter, 2018]
@@
 Learning Rate Schedule Cosine decay (no restarts) [Loshchilov and Hutter, 2017]
 Gradient Clipping Clipping above $\|\nabla\|=1$ [Rae et al., 2021]
@@
All models are parameterized by standard, decoder-only causally masked transformers with GeLU non-linearities [Hendrycks and Gimpel, 2016]. The models are implemented in Jax [Bradbury et al., 2018] and trained on Google TPU v3 hardware. We observe that large-scale pseudorandom number generation in Jax is slow, and therefore eschew the standard attention dropout regularization. The models are optimized using AdamW [Loshchilov and Hutter, 2018]. The learning rate schedule consists of a $1000$ step linear warmup to a maximum learning rate, followed by a single cycle of cosine decay [Loshchilov and Hutter, 2017] over the remaining steps to a final learning rate of zero. Following Chinchilla compute-optimality recommendations, we train each model for a number of steps that is approximately proportional to the model’s size [Hoffmann et al., 2022]. Configuration details of the models and optimization are presented in Table 6.
@@
Appendix G Details of the FIGARO Music Transformer Baseline
We compare our models to the implementation of Music Transformer [Huang et al., 2018] described by von Rütte et al. [2023]. We use the official public implementation for training and sampling from this model. [^4] https://github.com/dvruette/figaro While both our models and this baseline are based on the Transformer architecture and trained on Lakh MIDI, there are at several notable distinctions. The FIGARO Music Transformer architecture implements the relative attention mechanism proposed by [Huang et al., 2018]. The FIGARO Music Transformer inputs are tokenized using a REMI encoding that accounts for metrical structure [Huang and Yang, 2020]. The FIGARO Music Transformer is also a smaller-scale model than the other models considered in this paper: a six-layer transformer (approximately $30$M parameters) trained for $100$k iterations on sequences of length $256$. Some of these factors (e.g., model scale, sequence length) clearly favor the models presented in this paper; others (e.g., relative attention) may favor the FIGARO model. With so many uncontrolled variables, we caution against drawing conclusions about individual engineering choices in the design of the FIGARO Music Transformer versus the models proposed in this paper.
While pre-trained checkpoints of the FIGARO models are available, the training and evaluation splits used for the pre-trained checkpoints are incompatible with the splits defined in Section D. Therefore, we re-train our own version of the model using hashes 0d as the training split. A comparison of our re-trained model and the reference model checkpoint are presented in Figure 8. For definitions and discussion of the evaluation metrics compared here, see von Rütte et al. [2023]. Our version of the FIGARO model matches or slightly outperforms the reference model on these metrics.
@@ Appendix G Details of the FIGARO Music Transformer Baseline
Table 8: Quantitative evaluation metrics for the FIGARO implementation [von Rütte et al., 2023] of Music Transformer [Huang et al., 2018] trained on the FIGARO training data split, vs. the split defined in Section D.
@@
Appendix H A Baseline Autoregressive Infilling Algorithm
@@
$\mathbf{a}_{0}\leftarrow\texttt{SEP}$ ;
@@
       Sample $\mathbf{e}\sim p(\cdot|\mathbf{a}_{i-M:i-1})$ ;
@@
       $\mathbf{t}\leftarrow\text{Time}(\mathbf{e})$ ;
@@
       while $\emph{\text{Time}}(\mathbf{u}_{k})\leq\mathbf{t}$ ;
@@
             $\mathbf{a}_{i}\leftarrow\mathbf{u}_{k}$ ;
@@
      $\mathbf{a}_{i}\leftarrow\mathbf{e}$ ;
@@
Appendix I Details of Human Evaluation
@@
We compare accompaniments using (i) anticipatory autoregressive sampling (Algorithm 1) versus (ii) baseline autoregressive sampling. We also compare to (iii) baseline completions randomly sampled from the test set and (iv) skyline human compositions. When sampling without anticipation, we insert events from the melody into the conditional history of the model once generation passes them; we describe this modified sampling procedure formally in Appendix H. For each of $20$ three-bar prompts and single-part continuations, we generated accompaniments from each model. For each of the $12020\times\binom{4}{2}=$ 120 pairwise comparisons between continuations (as well as the baseline and skyline) we asked three human evaluators to indicate which clip is more musical, or that the clips are equally musical. Workers performed a total of $360$ comparisons.
@@
Medium (Row 8) FIGARO $95$ $24$ $31$ $1.187\times 10^{-08}$
@@
FIGARO Small (Row 3) $36$ $16$ $98$ $8.509\times 10^{-08}$
@@
Appendix J Training Optimization Logs
@@
Appendix K Model Card
Table 11: Model Card [Mitchell et al., 2019] - Anticipatory Music Transformer.
@@
Organization Developing the Model
Stanford Center for Research on Foundation Models
Model Date
June 2023
Model Type
Autoregressive Causal Transformer
Additional Modeling Details
See Section 3
License
Apache License, Version 2.0
Correspondence
jthickstun@cs.stanford.edu
@@
Primary Intended Uses
Collaborative co-composition between a human composer and an Anticipatory Music Transformer. The role of the anticipatory model in this collaboration could include, e.g., infilling tedious/low-entropy details (productivity enhancement) and suggesting possible continuations (creative ideation).
Primary Intended Users
Artists, musicians, and composers.
Out-of-Scope Uses
Long-Context Generation. These models cannot generate full-length song structures without human control. The models have a context length of $1024$ tokens ($331$ events). At $68$ tokens/second (the average for Lakh MIDI; see Appendix D) this corresponds to approximately 15 seconds of context. Models conditioned on more than $331$ events will only use the most recent $331$ events (including anticipated events) to predict the next event.
@@
Western Bias
These models are trained on the Lakh MIDI dataset, a collection of predominantly Western music. See Section 7.3 for further discussion.
@@
Automatic Metrics
Next-event perplexity (defined in Table 1) and bits per seconds (defined in Appendix E).
Human Evaluation
Pairwise human preferences between generated music and reference compositions.
Decision Thresholds
For human evaluation, we generated samples from anticipatory models using nucleus sampling with $p=0.95$. See Section 4 for further discussion.
Approaches to uncertainty and variability
We report p-values for pairwise comparisons between music generated by different models and ground truth music using the Wilcoxon signed-rank test. Due to computational constraints, we do not account for variability in the model training process, such as dataset splits or the random seed for optimization.
@@
Training Data
The 0d splits of the Lakh MIDI dataset, augmented using anticipation (see Section 3) with the prior distribution over controls described in Appendix C.
Validation Data
The e split of the Lakh MIDI dataset.
Test data
The f split of the Lakh MIDI dataset.
Out-of-Distribution Data
We do not evaluate out-of-distribution performance.
Preprocessing
Preprocessing and filtering of the Lakh MIDI dataset is described in Appendix D.
Motivation
We chose to work with the Lakh MIDI dataset because it is the largest collection of symbolic music data currently in use by the machine learning community.
@@
Aggregated Analysis
Our analysis of aggregate results based on automatic metrics and human evaluation are presented in Section 4.1 and Section 4.2 respectively. Key findings include: Anticipatory training does not interfere with autoregressive model performance, as measured by perplexities of comparable anticipatory and autoregressive models. Accompaniments generated by an Anticipatory Music Transformer have similar musicality to ground truth accompaniments according to human evaluators.
Disaggregated Analysis
We do not perform a disaggregated analysis of the Anticipatory Music Transformer. One obstruction to conducting such an analysis is a lack of metadata associated with the Lakh MIDI dataset.
@@
Labor Displacement
We are broadly concerned by the transient disruptions of labor markets caused by the introduction of new productivity-enhancing and automative technologies. See Section 7.1 for a discussion of the possible disruptive effects of generative music models on the creative economy.
Copyright
The Lakh MIDI dataset contains large quantities of copyrighted music. The copyright status of models trained on this data—and music sampled from these models—is an open legal question. See Section 7.2 for further discussion.

### v2

@@ Appendix A Ethical Considerations for Generative Models of Music
We find it likely that generative models of music will become widely deployed and used within the next several years (if not sooner). We are particularly concerned about the economic implications of this technology: how will these models affect labor markets for creative work (Section A.1)? We are uncertain about the legal status of these models: how will music created using models be treated by intellectual property and copyright law (Section A.2)? We are also curious how the advent of these models will change our perceptions and experience of art and creativity. Will widespread deployment of models trained on Western music reinforce existing Western hegemony in music culture (Section A.3)? Could the use of these models lead to creative stagnation, or loss of interest in music as an art form (Section A.4)? Analogous questions and concerns have been raised in the broader context of research on generative models, but the focus of this discussion is often grounded in generation of language or images. These discussions are worth revisiting in the specific context of music generation.
A.1 The Creative Economy and Labor Markets
@@
Music composition has economic value. What happens to the music economy—and to the people who rely on it for their livelihoods—if its creative product can be automatically generated by a machine? Economic research suggests that new technologies exert two countervailing effects on the labor market: a displacement effect whereby automation reduces demand for labor, and a productivity effect whereby greater productivity increases demand for labor (Acemoglu & Restrepo, 2018). It is unclear to us what the net effect of these two forces will be on the size of the labor market for music composition. If latent demand for new music is high, then decreasing the cost of music production could lead to more employment opportunities for composers. That said, this reconfiguration of the music economy might be disruptive, and success in the new music economy could require a new skillset that is difficult (or unappealing) for current workers to acquire. We suspect that generative music technology will affect the current music economy, and may displace some workers.
@@ A.2 Copyright
Regarding the outputs of a generative music model: generative models have been shown to plagiarize training data in both the language (Carlini et al., 2023) and vision (Somepalli et al., 2023) domains. Based on this evidence, we are certain that generative models of music are similarly capable of plagiarism. Technical mitigation strategies to prevent this behavior based on differential privacy or near access-freeness are nascent (Vyas et al., 2023), as are socio-technical strategies, such as providing copyright holders an opt-out mechanism to exclude their content from training datasets (Henderson et al., 2023). The models trained in this paper provide no technical guarantees against plagiarism. Therefore, any outputs of this model should be presumed to risk infringing on copyrighted material.
@@
A.3 Western Bias
While the methods developed in this paper are broadly applicable to modeling diverse musical traditions, the MIDI data format and the Lakh MIDI dataset used to train the anticipatory music transformer both impose a strong bias towards the Western music tradition. In Section 2 we adopted a musical vocabulary that describes musical pitch according to the 12-tone chromatic scale, which we inherit from the MIDI format of our training data. Models trained using this vocabulary are incapable of expressing music outside the 12-tone scale. This excludes, for example, the Koron and Sori quarter tones of Persian Dastgah music (Abdoli, 2011), the 53 Koma of Turkish makam (Benetos & Holzapfel, 2013), and the continuous-pitch gamaka appearing in Carnatic music (Viraraghavan et al., 2018). For further discussion of the consequences of modeling using a 12-tone discrete scale, see Lenchitz (2021).
@@ A.4 Models and Art
A.5 Releasing the Anticipatory Music Transformer
@@
Appendix B Licensing
@@
Appendix C Encoding Details for Anticipatory Infilling Models.
The tokenization of training examples (Section 3.3) for anticipatory infilling models using an arrival-time encoding of events (Definition 2.2) is described by Definition C.1. As discussed in Section 3.5, we double the base vocabulary to distinguish between anticipated events and non-anticipated events. We also include a sequence separator token and global control codes $\mathbf{z}\in\{\texttt{AR},\texttt{AAR}\}$ (Section 3.3) as well as the REST token (Section 3.2).
Definition C.1.
@@ 00 $10,048$ $11,060$ 00s $480$ms C4 piano
Example C.2.
See Table 4 for a structured description of this sequence.
The tokenization of training sequences for autoregressive models using an interarrival-time encoding (Defintion 2.3) is described by Definition C.3. We double the base vocabulary of notes and instruments to distinguish between onsets and offsets. We truncate interarrival times longer than 10 seconds (to 10 seconds).
Definition C.3.
[…]
If $\mathbf{x}^{\prime}_{i}$ is an offset,
[…]
And regardless,
[…]
Example C.4.
Appendix D A Prior over Music Infilling Controls
@@
Augmenting by a factor of $30$ results in an arrival-time encoded Lakh MIDI train set of approximately $52$B tokens. For training examples of length $1024$ and batch size $512$, this results in approximately one epoch per $100,000$ optimization steps. For models trained with larger step counts, we take multiple passes over the augmented dataset. Given the relatively small size of the Lakh MIDI dataset, anticipation may be of interest as a regularization technique (see Appendix K for some evidence of possible regularizing effects attributable to anticipation). In this case, it could be fruitful to (i) increase the augmentation factor further during training, and (ii) place a smaller weight on training without anticipation, to minimize the amount of duplication of the original (unaugmented) dataset within the augmented dataset.
Appendix E Details of the Lakh MIDI Dataset
@@
Figure 3: The distribution of instantaneous tokens/second calculated for the arrival-time tokenized Lakh MIDI validation split. Mean instantaneous tokens/second for the Lakh MIDI dataset is $68$ with a standard deviation of $51$ tokens/second.
Appendix F Measuring Cross-Entropy in Bits per Second
Because the arrival-time tokenization (Definition 2.2) and the interarrival-time tokenization (Defintion 2.3) describe nearly equivalent information, the log-loss of models trained using either encoding can be meaningfully compared via a unit conversion. To make comparison agnostic to encoding, we report losses in units of bits per second: this is the total log-loss of the test set, divided by the number of seconds of music in the test set. Concretely, given a per-token loss $L$ reported in nats per token, conversions for interrarival time and arrival time encodings of the test set defined in Section E are:
@@ Appendix F Measuring Cross-Entropy in Bits per Second
Truncations during tokenization (see Section 2.1) result in slightly different information content between an arrival-time tokenized sequence and the corresponding interarrival-time tokenized sequence. Under arrival-time tokenization, $0.05$% of note durations in the Lakh MIDI dataset exceed $10$ seconds, and are truncated (to $10$ seconds). Under interarrival-time tokenization, a total of $3947$ interarrival times ($0.00$%) in the Lakh MIDI dataset exceed $10$ seconds, and are truncated (to $10$ seconds). Finally, $0.09$% of training examples under arrival-time tokenization exceed $100$ seconds in length, and are discarded from the tokenized dataset. These instances are sufficiently rare that they exert a negligible effect on the information content (and therefore on bits per second comparisons) between the arrival-time and interarrival-time tokenized sequences.
Appendix G Hyperparameters and Resources
@@
       Optimizer       AdamW (Loshchilov & Hutter, 2018)
@@
       Learning Rate Schedule       Cosine decay (no restarts) (Loshchilov & Hutter, 2017)
       Gradient Clipping       Clipping above $\|\nabla\|=1$ (Rae et al., 2021)
@@
All models are parameterized by standard, decoder-only causally masked transformers with GeLU non-linearities (Hendrycks & Gimpel, 2016). The models are implemented in Jax (Bradbury et al., 2018) and trained on Google TPU v3 hardware. We observe that large-scale pseudorandom number generation in Jax is slow, and therefore eschew the standard attention dropout regularization. The models are optimized using AdamW (Loshchilov & Hutter, 2018). The learning rate schedule consists of a $1000$ step linear warmup to a maximum learning rate, followed by a single cycle of cosine decay (Loshchilov & Hutter, 2017) over the remaining steps to a final learning rate of zero. Following Chinchilla compute-optimality recommendations, we train each model for a number of steps that is approximately proportional to the model’s size (Hoffmann et al., 2022). Configuration details of the models and optimization are presented in Table 6.
@@
Appendix H Details of the FIGARO Music Transformer Baseline
We compare our models to the implementation of Music Transformer (Huang et al., 2018) described by von Rütte et al. (2023). We use the official public implementation for training and sampling from this model. [^4] https://github.com/dvruette/figaro While both our models and this baseline are based on the Transformer architecture and trained on Lakh MIDI, there are at several notable distinctions. The FIGARO Music Transformer architecture implements the relative attention mechanism proposed by (Huang et al., 2018). The FIGARO Music Transformer inputs are tokenized using a REMI encoding that accounts for metrical structure (Huang & Yang, 2020). The FIGARO Music Transformer is also a smaller-scale model than the other models considered in this paper: a six-layer transformer (approximately $30$M parameters) trained for $100$k iterations on sequences of length $256$. Some of these factors (e.g., model scale, sequence length) clearly favor the models presented in this paper; others (e.g., relative attention) may favor the FIGARO model. With so many uncontrolled variables, we caution against drawing conclusions about individual engineering choices in the design of the FIGARO Music Transformer versus the models proposed in this paper.
While pre-trained checkpoints of the FIGARO models are available, the training and evaluation splits used for the pre-trained checkpoints are incompatible with the splits defined in Section E. Therefore, we re-train our own version of the model using hashes 0d as the training split. A comparison of our re-trained model and the reference model checkpoint are presented in Table 8. For definitions and discussion of the evaluation metrics compared here, see von Rütte et al. (2023). Our version of the FIGARO model matches or slightly outperforms the reference model on these metrics.
@@ Appendix H Details of the FIGARO Music Transformer Baseline
In Table 9, we report the quantitative metrics proposed by von Rütte et al. (2023) using 800 20-second continuations of 3-bar prompts, for both the FIGARO Music Transformer, and our Medium Anticipatory Music Transformer (Row 8 in Table 1). While the Anticipatory Music Transformer marginally outperforms the FIGARO Music Transformer in all categories, it is not clear that we can draw a strong conclusion from these results. The reported metrics were designed to measure reconstruction quality: how faithfully does a generated output reproduce the original? These metrics make less sense for evaluating open-ended generation: we expect that generated continuations would be different from the originals. One hypothesis for the consistent out-performance of the Anticipatory Music Transformer is that music is often self-similar, and perhaps this model is better able to capture these self-similarities of music.
Table 8: Quantitative evaluation metrics for the FIGARO implementation (von Rütte et al., 2023) of Music Transformer (Huang et al., 2018) trained on the FIGARO training data split, compared to the same model trained using the split defined in Section E.
@@
Table 9: Quantitative evaluation metrics for the FIGARO implementation (von Rütte et al., 2023) of Music Transformer (Huang et al., 2018) compared to a Medium Anticipatory Model (Row 8 in Table 1).
Model I $\uparrow$ C $\uparrow$ ND $\downarrow$ P $\uparrow$ D $\uparrow$ $s_{c}$uparrow $s_{g}$uparrow
von Rütte et al. (2023) $0.833$ $0.364$ $0.574$ $0.751$ $0.573$ $0.619$ $0.257$
Anticipatory $0.836$ $0.392$ $0.529$ $0.755$ $0.593$ $0.659$ $0.259$
Appendix I A Baseline Autoregressive Infilling Algorithm
@@
$\mathbf{a}_{0}\leftarrow\texttt{SEP}$
@@
       Sample $\mathbf{e}\sim p(\cdot|\mathbf{a}_{i-M:i-1})$
@@
       $\mathbf{t}\leftarrow\text{Time}(\mathbf{e})$
@@
       while $\emph{\text{Time}}(\mathbf{u}_{k})\leq\mathbf{t}$
@@
             $\mathbf{a}_{i}\leftarrow\mathbf{u}_{k}$
@@
      $\mathbf{a}_{i}\leftarrow\mathbf{e}$
@@
Appendix J Details of Human Evaluation
@@
We compare accompaniments using (i) anticipatory autoregressive sampling (Algorithm 1) versus (ii) baseline autoregressive sampling. We also compare to (iii) baseline completions randomly sampled from the test set and (iv) skyline human compositions. When sampling without anticipation, we insert events from the melody into the conditional history of the model once generation passes them; we describe this modified sampling procedure formally in Appendix I. For each of $20$ three-bar prompts and single-part continuations, we generated accompaniments from each model. For each of the $12020\times\binom{4}{2}=$ 120 pairwise comparisons between continuations (as well as the baseline and skyline) we asked three human evaluators to indicate which clip is more musical, or that the clips are equally musical. Workers performed a total of $360$ comparisons.
@@
Medium (Row 8) Music Transformer $95$ $24$ $31$ $1.187\times 10^{-08}$
@@
Music Transformer Small (Row 3) $36$ $16$ $98$ $8.509\times 10^{-08}$
@@
Appendix K Training Optimization Logs
@@
Appendix L Model Card
Table 12: Model Card (Mitchell et al., 2019) - Anticipatory Music Transformer.
@@
Organization Developing the Model Stanford Center for Research on Foundation Models
Model Date June 2023
Model Type Autoregressive Causal Transformer
Additional Modeling Details See Section 3
License Apache License, Version 2.0
Correspondence jthickstun@cs.stanford.edu
@@
Primary Intended Uses Collaborative co-composition between a human composer and an Anticipatory Music Transformer. The role of the anticipatory model in this collaboration could include, e.g., infilling tedious/low-entropy details (productivity enhancement) and suggesting possible continuations (creative ideation).
Primary Intended Users Artists, musicians, and composers.
Out-of-Scope Uses Long-Context Generation. These models cannot generate full-length song structures without human control. The models have a context length of $1024$ tokens ($331$ events). At $68$ tokens/second (the average for Lakh MIDI; see Appendix E) this corresponds to approximately 15 seconds of context. Models conditioned on more than $331$ events will only use the most recent $331$ events (including anticipated events) to predict the next event.
@@
Western Bias These models are trained on the Lakh MIDI dataset, a collection of predominantly Western music. See Section A.3 for further discussion.
@@
Automatic Metrics Next-event perplexity (defined in Table 1) and bits per seconds (defined in Appendix F).
Human Evaluation Pairwise human preferences between generated music and reference compositions.
Decision Thresholds For human evaluation, we generated samples from anticipatory models using nucleus sampling with $p=0.95$. See Section 4 for further discussion.
Approaches to uncertainty and variability We report p-values for pairwise comparisons between music generated by different models and ground truth music using the Wilcoxon signed-rank test. Due to computational constraints, we do not account for variability in the model training process, such as dataset splits or the random seed for optimization.
@@
Training Data The 0d splits of the Lakh MIDI dataset, augmented using anticipation (see Section 3) with the prior distribution over controls described in Appendix D.
Validation Data The e split of the Lakh MIDI dataset.
Test data The f split of the Lakh MIDI dataset.
Out-of-Distribution Data We do not evaluate out-of-distribution performance.
Preprocessing Preprocessing and filtering of the Lakh MIDI dataset is described in Appendix E.
Motivation We chose to work with the Lakh MIDI dataset because it is the largest collection of symbolic music data currently in use by the machine learning community.
@@
Aggregated Analysis Our analysis of aggregate results based on automatic metrics and human evaluation are presented in Section 4.1 and Section 4.2 respectively. Key findings include: Anticipatory training does not interfere with autoregressive model performance, as measured by perplexities of comparable anticipatory and autoregressive models. Accompaniments generated by an Anticipatory Music Transformer have similar musicality to ground truth accompaniments according to human evaluators.
Disaggregated Analysis We do not perform a disaggregated analysis of the Anticipatory Music Transformer. One obstruction to conducting such an analysis is a lack of metadata associated with the Lakh MIDI dataset.
@@
Labor Displacement We are broadly concerned by the transient disruptions of labor markets caused by the introduction of new productivity-enhancing and automative technologies. See Section A.1 for a discussion of the possible disruptive effects of generative music models on the creative economy.
Copyright The Lakh MIDI dataset contains large quantities of copyrighted music. The copyright status of models trained on this data—and music sampled from these models—is an open legal question. See Section A.2 for further discussion.